OpenStudy (anonymous):

The area of a rectangle is given by A = x2 + 4x – 5. Find the value of x if the total area of the rectangle is equal to 7 square units

OpenStudy (anonymous):

x2 + 4x – 5=7 x2 + 4x –12=0 (x+6)(x-2) x=-6 or 2